OpenStudy (anonymous):

Can anyone help me with this question? I have no idea how to approach getting the velocities: 4. A basketball leaves the player’s hands at a height of 2.10 m above the floor at a 35.0o angle to the floor. The shooter is standing 12.0 m from the basket and the rim of the basket is 2.60 m above the ground. If the rim is 22.0 cm across, what range of initial speeds does the shooter have to sink the ball?

OpenStudy (anonymous):

solve it in general for Vi, leaving the x distance as a parameter, then plug in your two different values of x. I guess they would be 12 and 12.22.... question is a bit vague on this point, since width of the ball matters and what distance exactly is measured by the '12m'. Assuming the '12m' is to the front of the rim, and any shot that hits there goes in then go with 12 and 12.22m .5 = Vi*sin35*t - 1/2 * g * t^2 d = Vi*cos35*t t=d/(Vi*cos35) subs into the y equation and solve for Vi